1. A very tall skycraper sways back and forth every 4.0 seconds.

Period:

Frequency:


2. A pendulum takes 0.5 second to complete one cycle.

Period:

Frequency:


1
Expert's answer
2021-12-22T09:25:20-0500

1) Period:


"T=4.0\ s."

Frequency:


"f=\dfrac{1}{T}=\dfrac{1}{4.0\ s}=0.25\ Hz."

2) Period:


"T=0.5\ s."

Frequency:


"f=\dfrac{1}{T}=\dfrac{1}{0.5\ s}=2\ Hz."